OpenStudy (anonymous):

5/6 = n/24 , The was that was given was wrong anyone ?

OpenStudy (anonymous):

n = 20

OpenStudy (anonymous):

cross product property. 24*5=6*n==>120=6*n so n=20

OpenStudy (anonymous):

Thank you :)